OpenStudy (ghdoru11):

Find the Fourier expansion for f(x)=4x, x[-pi,pi], T=2 pi

OpenStudy (3mar):

This is the expansion of f(x)=x You can multiply by 4 to meet yours.

OpenStudy (astrophysics):

.

OpenStudy (kevin):

@3mar is awesome, he can help you :D

OpenStudy (irishboy123):

f(x) = 4x is odd so you only need the sine terms of the expansion So coefficients are: \(b_n = \dfrac{1}{\pi}\int\limits_{-\pi}^{\pi} 4x \, \sin nx\, dx\) \( = \dfrac{8}{\pi}\int\limits_{0}^{\pi} x \dfrac{d}{dx}(- \dfrac{1}{n}\cos(nx) ) \, dx\) ie do it by IBP

OpenStudy (3mar):

Thanks for the medal! Hope I could help you.